0.202 gm of a carbon compound on combustion , gave 0.361 gm of carbon dioxide and 0.47 gm of water . calculate the percentage composition of carbon :

0.202 gm of  a carbon compound on combustion , gave 0.361 gm of carbon dioxide and 0.47 gm of water . calculate the percentage composition of carbon :

Grade:9

1 Answers

Shivangi Khatter
5 years ago
moles of CO2= 0.361/44 = 0.0082
mass of carbon = 0.0082 * 12 = 0.095845
percentage of carbon = 0.09845/0.202 * 100 = 48.73 %